OpenStudy (anonymous):

Gannon has deposited $742 in a savings account that earns interest at a rate of 3.4% compounded monthly. What will the account balance be in 15 years?

OpenStudy (anonymous):

\[742\times (1+\frac{.034}{12})^{12\times 15}\] and a calculator

OpenStudy (anonymous):

i get 1234 http://www.wolframalpha.com/input/?i=742 \times+%281%2B\frac{.034}{12}%29^{12\times+15}

OpenStudy (anonymous):

my calcualtor says error